OpenStudy (anonymous):

PlEaSe HeLp Tell whether the lines for each pair of equations are parallel, perpendicular or neither. 1) y=-2x+4 and -5x+10y=5 2) y=(-1/4)x+8 and -2x+8y=4 Thanks so much!

OpenStudy (anonymous):

How do you know when a set of lines are parallel, or perpendicular?

OpenStudy (anonymous):

y=-2x+4 and y = 1/2x+1/2 y=(-1/4)x+8 and y= 1/4x+1/2 parallel has the same slope perpendicular has a slope that is a reciprocal with the opposite sign

OpenStudy (anonymous):

in this case, which one is has the same slope, or reciprocal with opposite sign or neither